# Differentially Private Matchings

### v2

@@ 5.2.2 Utility and Number of Rounds
For any $\varepsilon\in(0,1)$, Algorithm 3 returns a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum matching, with high probability, when $b=\Omega(\log^{2}(n)/\varepsilon)$, in $O(\log(n))$ rounds.
Towards proving Lemma 5.2, we first prove the following lemma. Recall that our algorithm randomly divides the population into (roughly) half proposers and half receivers. Then, the proposers first propose their proposal sets, and then the receivers respond with their match sets for each of $32c\cdot\lceil\log_{16/15}(n)\rceil$ rounds. The following crucial lemma gives the “progress” of the algorithm after each round. Namely, to ensure we produce a maximal matching in $O(\log n)$ rounds, with high probability, roughly a constant fraction of the nodes would need to exceed their matching capacity every round (hence, would not participate in future matching rounds). For convenience in the below proofs, when we write $\log(n)$, we mean $\log_{1+\eta}(n)$.
In the analysis below, we call the set of nodes that have not exceeded their matching capacity and have a non-zero number of neighbors who have not reached their matching capacity, the hopeful nodes. A node that is not hopeful is called unhopeful.
First, note that every active node in any round $i$ is either a proposer or receiver by definition. We first prove the following lemma, which states that a vertex $w$ will choose $r_{w}=0$ with high probability, as a proposer or receiver if its degree is less than $27c\log(n)/\varepsilon$.
@@
Suppose $X\sim\text{Lap}(4/\varepsilon^{\prime})$ and $Y\sim\text{Lap}(2/\varepsilon^{\prime})$ are the noise random variables chosen in Algorithms 4 and 4, respectively. Since $w$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$, it will pick $r_{w}=0$ if $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$. By Lemma A.1, $X\leq 24c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$ and $Y\leq 12c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$. Then, $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{5c}}$ when $n\geq 2$. ∎
We now prove an additional lemma about the size of any proposal or match set an active vertex in round $i$ will choose. In particular, if a vertex is active in round $i$, then it will propose a proposal set or match set of size at least $32c\log_{16/15}(n)/\varepsilon^{\prime}$, with high probability.
@@
If vertex $w\in V_{i}$ is active in round $i$, then $w$ will pick $r_{w}=0$ or an $r_{w}$ where $|W_{r_{w}}(w)|\geq 32c\log_{16/15}(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{2c}}$.
@@
If $w$ is still active in round $i$, this means that $w$ did not meet its matching condition in round $i-1$. By Lemma A.1, this means that with probability at least $1-\frac{1}{n^{4c}}$, it holds that $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ because the noise drawn in Algorithm 3 is at least $-12c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$ and the noise drawn in Algorithm 3 is at most $24c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since, $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}<<27c\log(N)/\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}>>\varepsilon^{\prime}$, by our proof for Lemma 5.3, node $w$ will choose a $|W_{r_{w}}(w)|$ can be as large as $50c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since we chose $\eta=0.5$, we will pick an $r_{w}$ where $33c\log(n)/\varepsilon^{\prime}<(50\cdot 2)/3c\log(n)/\varepsilon^{\prime}\leq\mathbb{E}[|W_{r}(w)|]\leq 50c\log(n)/\varepsilon^{\prime}$. Otherwise, if the expectation is less than $33c\log(n)/\varepsilon^{\prime}$, then $w$ will pick all available edges (and hence pick $r_{w}=0$). Now we show the concentration of the size with sufficiently high probability via the Chernoff bound. We can use the Chernoff bound since the edges picked for $W_{r_{w}}(n)$ are i.i.d. chosen at random using coins that are flipped heads with probability $(1+\eta)^{-r_{w}}$. Specifically, by Theorem A.2, if we take $\psi=0.4$, then, $|W_{r_{w}}(w)|<32c\log_{16/15}(n)/\varepsilon^{\prime}$ with probability at most $\exp\left(-\frac{0.4^{2}\cdot 33c\log(n)}{3\varepsilon^{\prime}}\right)<\frac{1}{n^{3c}}$. Hence, together, $w$ will pick an $r_{w}$ where $|W_{r_{w}}(w)|\geq 32c\log_{16/15}(n)/\varepsilon^{\prime}$ or it will pick $r_{w}=0$. ∎
@@
By definition of a hopeful edge, $e=\{u,v\}$, if either $u$ or $v$ becomes a proposer then they will propose a non-empty set (if they have at least one receiver neighbor). The marginal probability that any hopeful edge has one endpoint become the proposer and the other remains a receiver is $1/4$ by Algorithm 3. We now follow a charging scheme inspired by the edge orientation charging scheme given in [khuller1994primal]; however, our charging scheme is fundamentally different since we are performing charging in a noisy process for $b$-matching. Note that we do not actually perform any orientation algorithm; this orientation charging scheme is only used for the analysis. In this orientation charging scheme, we orient all hopeful edges from low to high degree endpoints in the induced graph consisting of $H$. We denote a vertex $v$ as good if at least $1/3$ of its degree is oriented into it. We call an edge good if it is oriented into a good vertex. We first show that at least $1/2$ of $H$ is good. We call a vertex or edge bad if it is not good.
A vertex is bad if more than $2/3$ of its degree is oriented out. These edges that are oriented out are bad edges if and only if they are oriented toward bad vertices. Furthermore, any bad vertex must either have degree $1$ or have at least $2$ outgoing edges for every bad incoming edge. A degree one vertex has a bad outgoing edge if it is paired with another vertex with at least two outgoing edges. Hence, for any node we can charge each incoming bad edge to a unique outgoing edge such that each outgoing edge receives at most one charge. Suppose the number of incoming bad edges charged to good edges is $H^{\prime}$, then there are at least $H^{\prime}$ good edges. Finally, there can be at most $\frac{1}{2}(H-H^{\prime})$ bad edges charged to bad edges, since any incoming bad edge must be charged to at most one outgoing edge and every bad vertex has at least two outgoing edges for each incoming edge. Hence, in total, at least $H/2$ edges are good edges.
We now show that at least a fixed constant fraction of the hopeful, good edges become unhopeful each round with high constant probability. For each good, hopeful edge $e=(u,v)$, oriented from $u$ to $v$, the marginal probability that the source of the edge is a proposer and the sink is a receiver is $1/4$. Let the degree of $v$ in the induced subgraph consisting of all hopeful edges be $\deg(v)$; since $e$ is oriented from $u$ to $v$, it means that $\deg(u)\leq\deg(v)$. The marginal probability that $u$’s proposed set contains $v$ is at least
$\min\left(1,\frac{32c\log_{16/15}(n)}{\varepsilon^{\prime}\cdot\deg(u)}\right)\geq\min\left(1,\frac{32c\log_{16/15}(n)}{\varepsilon^{\prime}\cdot\deg(v)}\right)$
by Lemma 5.4. Thus, the probability that $u$ is a proposer, $v$ is a receiver, and $u$ proposes $v$ is at least $\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$. Receiver $v$ can either choose to match with $u$ or not. If $v$ does not match with $u$, that means by Lemma 5.4, $v$ picked a set that did not contain $u$ and has reached its matching threshold, with probability at least $1-\frac{1}{n^{3c}}$.
The expected number of edges $v$ receives as part of all proposals is then
$\deg(v)\cdot\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(u)}\right)$ $\geq\deg(v)\cdot\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$
$=\min\left(\deg(v),\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}}\right)\,.$
Thus, in expectation, $v$ satisfies its matching condition. Now, we show the concentration on the probability that $v$ satisfies the matching condition. By the Chernoff bound and suppose $\psi=0.4$, the probability that $<10c\log(n)/(4\varepsilon^{\prime})$ proposers propose to $v$ is upper bounded by
$\exp\left(\frac{0.4^{2}\cdot 32c\log_{16/15}(n)/\varepsilon^{\prime}}{12}\right)<\exp\left(\frac{2c\log(n)}{12\varepsilon^{\prime}}\right)<\frac{1}{n^{4c}}\,.$
Hence, by the union bound over $n$ nodes, with probability at least $1-\frac{1}{n^{3c}}$, each good receiver satisfies its matching condition. Each hopeful, good edge has $1/4$ probability that it is going from a proposer to a receiver. The expected number of hopeful edges that are good and are going from proposers to receivers is then at least $\frac{H}{8}\cdot(1-1/n^{3})$; these are also the expected number of hopeful edges that become unhopeful. By the Markov bound, the probability that at most $1615H/$ 16 hopeful edges remain hopeful is $\frac{7/8+\frac{1}{n^{3}}}{15/16}=14/15+\frac{15}{16n^{3}}$. This means that with probability at least $1/16$ (assuming $n>10$—if $n\leq 10$, we can run our algorithm from Section 4 in $O(1)$ rounds), at least $H/16$ of the hopeful edges become unhopeful. ∎
@@
By Lemma 5.5, in each round $i$, at least $1/16$ of the hopeful edges become unhopeful with probability at least $1/16$. In $32\cdot 16c\log(n)$ rounds, the expected number of rounds for which at least $1/16$ of the hopeful edges become unhopeful is $32c\log(n)$ rounds. Because the Chernoff bound holds for binomials, by Theorem A.2, if we set $\psi=0.5$, then the probability that less than $16c\log(n)$ rounds are successful is at most $\exp\left(\frac{0.5^{2}\cdot 32c\log(n)}{3}\right)<\frac{1}{n^{8c}}$. Hence, in $O(\log n)$ rounds, all edges become unhopeful.
We now show that if all edges become unhopeful, then all nodes are either matched or all of their neighbors are matched. An edge becomes unhopeful if at least one of its endpoints satisfies the matching condition. By Lemma A.1, a node satisfies the matching condition if and only if $|M(u)|\geq 1$ with probability at least $1-\frac{1}{n^{2c}}$. Hence, if a node is adjacent to all unhopeful edges, then it either satisfies the matching condition and is matched to at least one neighbor with high probability, or it is adjacent to endpoints which are matched with high probability.
Finally, by Lemma A.1, $|M(u)|\geq 2\cdot 147c\log(n)/\varepsilon^{\prime}$ with probability at most $\frac{1}{n^{2c}}$. By Lemma 4.2 and a union bound over all probabilities of failure, with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$, we output a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum matching in $O(\log n)$ rounds when $b=\Omega(\log^{2}(n)/\varepsilon)$.
The number of rounds of our algorithm is determined by Algorithm 3 (which is $O(\log n)$) since round contains a constant number of synchronization points and all nodes (proposers and receivers) perform their instructions simultaneously. ∎
@@
6 Node Differentially Private Implicit Matchings
Using sparsification techniques, we demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. In particular, we prove Theorem 1.8, which we restate below. See 1.8
The class of bounded arboricity [^8] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree $n-1$ and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. For simplicity of presentation, we first design an algorithm assuming a public bound $\tilde{\alpha}$. Then our algorithm is always private but the approximation guarantees hold when $\tilde{\alpha}$ upper bounds the arboricity of the input graph. When such a public bound is unavailable, we show that guessing the bound with powers of 2 achieves the same guarantees up to logarithmic factors.
@@ 6 Node Differentially Private Implicit Matchings
6.1 Bounded Arboricity Sparsifiers
[…]
Matching Sparsifier
[…]

### v3

@@ 6.2.2 Utility and Number of Rounds
For any $\varepsilon\in(0,1)$, Algorithm 3 returns a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum $b^{\prime}$-matching, with high probability, when $b$ satisfies the lower bound in Theorem 2.3, in $O(\log(n))$ rounds.
Towards proving Lemma 6.4, we first prove the following lemma. Recall that our algorithm randomly divides the population into (roughly) half proposers and half receivers. Then, the proposers first propose their proposal sets, and then the receivers respond with their match sets for each of $32c\cdot\lceil\log_{16/15}(n)\rceil$ rounds. The following crucial lemma gives the “progress” of the algorithm after each round. Namely, to ensure we produce a maximal $b^{\prime}$-matching in $O(\log n)$ rounds, with high probability, roughly a constant fraction of the nodes would need to exceed their $b^{\prime}$-matching condition every round; hence, these nodes would not participate in future matching rounds. For convenience in the below proofs, when we write $\log(n)$, we mean $\log_{1+\eta}(n)$.
In the analysis below, we call the set of nodes that have not exceeded their $b^{\prime}$-matching condition and have a non-zero number of neighbors who have not reached their $b^{\prime}$-matching condition, the hopeful nodes. A node that is not hopeful is called unhopeful.
First, note that every active node in any round $i$ is either a proposer or receiver by definition. We first prove the following lemma, which states that a vertex $w$ will choose $r_{w}=0$ with high probability, as a proposer or receiver if its degree is less than $27c\log(n)/\varepsilon^{\prime}$.
@@
Suppose $X\sim\text{Lap}(4/\varepsilon^{\prime})$ and $Y\sim\text{Lap}(2/\varepsilon^{\prime})$ are the noise random variables chosen in Algorithms 4 and 4, respectively. Since $w$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$, it will pick $r_{w}=0$ if $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ by our setting of $b$ in Algorithm 3 of Algorithm 3 and Algorithm 4 of Algorithm 4. By Lemma A.1, $X\leq 24c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$ and $Y\leq 12c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$. Then, $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{5c}}$ when $n\geq 2$. ∎
We now prove an additional lemma about the size of any proposal or match set an active vertex in round $i$ will choose. In particular, if a vertex is active in round $i$, then it will propose a proposal set or match set of size at least $32c\log(n)/\varepsilon^{\prime}$, with high probability.
@@
If vertex $w\in V_{i}$ is active in round $i$, then $w$ will pick $r_{w}=0$ or an $r_{w}$ where $|W_{r_{w}}(w)|\geq 20c\log(n)/\varepsilon^{\prime}$ with probability at least $1-O\left(\frac{1}{n^{2c}}\right)$.
@@
If $w$ is still active in round $i$, this means that $w$ did not meet its $b^{\prime}$-matching condition in round i1i-1. Hence, by Algorithm 3 of Algorithm 3 and Lemma A.1, it holds that $|M(w)|<b-259c\log(n)/\varepsilon^{\prime}+\text{Lap}(4/\varepsilon^{\prime\prime})-\text{Lap}(8/\varepsilon^{\prime\prime})$. By Lemma A.1 and the setting of $\varepsilon^{\prime\prime}$, we can bound $b-439c\log(n)/\varepsilon^{\prime}\leq|M(w)|\leq b-79c\log(n)/\varepsilon^{\prime}$ with probability at least $1-O\left(\frac{1}{n^{2c}}\right)$. Since our proposal/match selection guarantees $|\widetilde{M}(w)|+|\widetilde{W}_{r}(w)|+27c\log(n)/\varepsilon^{\prime}\leq b$, there is sufficient slack to choose an index $r$ with $\mathbb{E}[|W_{r}(w)|]=40c\log(n)/\varepsilon^{\prime}$. By a Chernoff bound and independence of the public coins, with probability at least $1-\frac{1}{n^{5c}}$ we have $|W_{r_{w}}(w)|\geq 20c\log(n)/\varepsilon^{\prime}$, or else $r_{w}=0$ if the expectation is smaller (in which case all available edges are taken). ∎
@@
By definition of a hopeful edge, $e=\{u,v\}$, if either $u$ or $v$ becomes a proposer then they will propose a non-empty set (if they have at least one receiver neighbor). The marginal probability that any hopeful edge has one endpoint become the proposer and the other remains a receiver is $1/4$ by Algorithm 3. We use an orientation charging scheme as follows: in the induced graph on the endpoints of $H$, orient all hopeful edges from lower to higher degree. A vertex is good if at least $1/3$ of its incident hopeful edges are oriented into it; an oriented edge into a good vertex is good. A standard charging shows at least $H/2$ edges are good.
Fix a good, hopeful edge $e=(u,v)$ oriented from $u$ to $v$. The marginal probability that $u$ is a proposer and $v$ is a receiver is $1/4$. By Lemma 6.6, the marginal probability that $u$ proposes $v$ is at least $\min\!\left(1,\frac{20c\log(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$. The expected number of proposals received by $v$ is therefore at least $\min\!\left(\deg(v),\frac{20c\log(n)}{4\varepsilon^{\prime}}\right)$, so by a Chernoff bound, with probability at least $1-\frac{1}{n^{2c}}$ the receiver $v$ either accepts one of these proposals or reaches its $b^{\prime}$-matching condition in this round. Hence, with probability at least $1/4\cdot(1-1/n^{2c})$, a good, hopeful edge becomes unhopeful in this round. Taking expectations and applying Markov’s inequality as in the standard analysis yields the claim that at least $H/16$ edges become unhopeful with probability at least $1/16$ (for $n$ large enough). ∎
@@
By Lemma 6.7, in each round $i$, at least $1/16$ of the hopeful edges become unhopeful with probability at least $1/16$. In $32\cdot 16c\log(n)$ rounds, the expected number of rounds for which at least $1/16$ of the hopeful edges become unhopeful is $32c\log(n)$ rounds. By a Chernoff bound, the probability that fewer than $16c\log(n)$ rounds are successful is at most $n^{-4c}$. Hence, in $O(\log n)$ rounds, all edges become unhopeful.
We now show that if all edges become unhopeful, then all nodes are either matched to at least $b^{\prime}$ neighbors or all of their neighbors are matched to at least $b^{\prime}$ neighbors. An edge becomes unhopeful if at least one of its endpoints satisfies the $b^{\prime}$-matching condition. By Lemma A.1, a node satisfies the matching condition if and only if $|M(u)|\geq b^{\prime}$ with probability at least $1-\frac{1}{n^{2c}}$. Hence, if a node is adjacent to all unhopeful edges, then it either satisfies the matching condition and is matched to at least $b^{\prime}$ neighbors with high probability, or it is adjacent to endpoints which are matched with at least $b^{\prime}$ neighbors with high probability.
Finally, by Lemma A.1, $|M(u)|\geq b$ with probability at most $1/n^{2c}$, and this is ruled out by the choice of $r$ in Algorithm 4. By Lemma 6.6 and a union bound over all failure events, with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$, we output a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum $b^{\prime}$-matching in $O(\log n)$ rounds provided that $b$ satisfies the lower bound in Theorem 2.3.
The number of rounds of our algorithm is determined by Algorithm 3 (which is $O(\log n)$) since each round contains a constant number of synchronization points and all nodes (proposers and receivers) perform their instructions simultaneously. ∎
@@
6.3 Corollaries
By choosing a fixed constant $\eta=\nicefrac{{1}}{{2}}$, benchmark parameter $b^{\prime}=1$, and matching degree $b=\Omega\left(\frac{\log^{2}(n)}{\varepsilon}\right)$, Theorem 2.3 immediately implies Corollary 6.1
In order to obtain a non-bicriteria approximation when $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{5}\varepsilon}\right)$, we use the same technique as the sequential edge-DP algorithm, but this time we execute the $O(\log n)$-round algorithm (Theorem 2.3) with $b^{\prime}=\frac{b}{1+O(\eta)}$ instead of the sequential algorithm. The proof is identical to that of Corollary 5.7. This yields a proof of Theorem 6.2.
7 Node Differential Privacy via Arboricity Sparsification
Node-DP Matchings.
Using sparsification techniques, we demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. In particular, we prove Theorem 2.6, which we restate below for convenience. See 2.6
As a special case, we have the following corollary
Corollary 7.1 (Node-DP Approximate Maximum Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$ in the billboard model for $b=O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta\varepsilon^{2}}\right)$ that has the size of a $(2+\eta)$-approximate maximum matching.
Node-DP Vertex Cover.
To demonstrate the applicability of our techniques, we also design an implicit node-DP vertex cover algorithm using arboricity sparsification. Its guarantees are summarized in Theorem 2.7, which we repeat below. See 2.7
The class of bounded arboricity [^8] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree n1n-1 and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. For simplicity of presentation, we first design an algorithm assuming a public bound $\tilde{\alpha}$. Then our algorithm is always private but the approximation guarantees hold when $\tilde{\alpha}$ upper bounds the arboricity of the input graph. When such a public bound is unavailable, we show that privately estimating $\alpha$ yields the same guarantees up to logarithmic factors.
@@ 7 Node Differential Privacy via Arboricity Sparsification
Interestingly, in Section 7.4, we show that actually releasing such a sparsifier is impossible under privacy constraints (Theorem 2.5). However, we can still use them as subroutines in our algorithms!
Organization.
The rest of the section is organized as follows. In Section 7.1, we tackle node-DP matchings as a warm-up before addressing general $b$-matching. This entails adapting a matching sparsifier to the DP setting in Section 7.1.1, allowing us to reduce node-DP matching to edge-DP matching for bounded arboricity graphs. We treat the more general case of $b$-matchings in Section 7.2. This requires extending our matching sparsifier to a $b$-matching sparsifier in Section 7.2.1, which may be of independent interest beyond our work. To further illustrate the power of arboricity sparsification, we further design a node-DP vertex cover algorithm in Section 7.3. This includes a vertex cover sparsifier in Section 7.3.1, which allows us to reduce node-DP vertex cover to edge-DP vertex cover, similar to matchings. Finally, we present lower bounds for releasing our sparsifiers in Section 7.4.
7.1 Node-DP Implicit Matching
In this section, we design a node-DP matching algorithm via arboricity sparsification.
7.1.1 Bounded Arboricity Matching Sparsifier
For our bounded arboricity graphs, we take inspiration from the bounded arboricity sparsifier of [solomon2021local]. A closely related line of work is that of edge degree constrained subgraphs (EDCS) [assadi2019coresets, bernstein2015fully, bernstein2016faster]. We modify the sparsifier from [solomon2021local] to show Proposition 7.2, which states that node-neighboring graphs have small edge edit distance post-sparsification.
Matching Sparsifier.
Our sparsification algorithm $\textsc{ContractionSparsify}_{\pi}$ proceeds as follows. Given an ordering $\pi\in P_{\binom{n}{2}}$ over unordered vertex pairs, a graph $G$, and a degree threshold $\Lambda$, each vertex $v$ marks the first $\min(\deg_{G}(v),\Lambda)$ incident edges with respect to $\pi$. Then, $H$ is obtained from $G$ by taking all vertices of $G$ as well as edges that were marked by both endpoints. In the central model, we can take $\pi$ to be the lexicographic ordering of edges $\{u,v\}$. Thus, we omit the subscript $\pi$ in the below analyses with the understanding that there is a fixed underlying ordering.